OpenStudy (anonymous):

Graph of the linear equation Y+2=1/5(x+1)

OpenStudy (praxer):

5y+10=x+1 5y=x-9 y=x/5-9/5 so, the slope is 1/5 and the intercept is -9/5.. can you plot it now........ :)

OpenStudy (anonymous):

Thank u so much

OpenStudy (praxer):

Welcome....... :)